# Interpretable Multi-Source Data Fusion Through Latent Variable Gaussian Process

### v1

@@ 6 Conclusion
AI and ML frameworks have had a strong impact in the science and engineering communities. On the other hand, with the ever-rising application of data-driven surrogates on multiple sources of data, whether it is gathered experimentally or computationally, not much attention has been paid to their quality, completeness and comprehensiveness of non-included and unknown underlying parameters involved in data generation. These parameters vary from differences in manufacturing process, testing technique, and more. Through the research done in this paper, an interpretable multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed and extensively studied. Through LVGP models, a categorical variable was introduced to represent sources of information during model training and predictions. A low-dimensinal latent space is then learned to incorporate the categorical source variable into multi-source data fusion. Through latent variable representation, the non-included and underlying unknown physical parameter variations among the different sources is captured and the differences between sources is quantified and introduced to the modelling, making the multi-source ML model interpretable and aware with respect to the differences and similarities between the information sources. The inclusion of the data sources not only leads to more accurate models but also enables the study of source specific response surfaces, material optimization, and sensitivity analysis. Compared to using single source data for ML, the proposed data fusion method using multi-source information can provide a better prediction for sparse-data problems by taking advantage of the information of correlations and relationships among different data sources. With the quantified uncertainty from LVGP for the specified information source, the approach can further guide adaptive sampling or active learning when allocating resources for data collection. Additionally, a novel dissimilarity metric is also defined through the latent variables of the LVGP models that can help understand the aggregated differences and similarities in the information sources. The proposed approach is applied on and analyzed through two mathematical (representative parabola and 2D Ackley function) and two materials design (design of FeCrAl alloys and design of SmCoFe alloys) case studies. Finally, a novel application of the framework would involve performing materials design and development through autonomous experimentation studies where multiple sources of experimental data are gathered to guide the study. It is envisioned that the proposed multi-source framework is not limited to experimental or computational information sources, and can be further extended to mixture of experimental and computational sources in numerous engineering and science applications.
@@ Acknowledgments
This work is supported by the US Department of Energy, National Nuclear Security Administration, under award number DE-NE0009047. Northwestern researchers are grateful for the support from NSF 2219489 and NSF 2037026. This material is based upon work supported by the U.S. Department of Energy’s Office of Energy Efficiency and Renewable Energy (EERE) under the Advanced Manufacturing Office, Award Number DE-AC0206H11357. The views expressed herein do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
@@
Information Source
Function Formulation $(a=20,b=0.2,c=2)$
Number of Training Samples
Number of Testing Samples
Ground Source
$z=-a*exp(-b(\sqrt{\frac{1}{2}(x^{2}-y^{2})}))-exp(\frac{1}{2}(cos(cx)+cos(cy)))+a+exp(1)$
$20$
$100$
Perturbed Source - 1
$-a*exp(-b(\sqrt{\frac{1}{2}(x^{2}-y^{2})}))+10$
$50$
$100$
Perturbed Source - 2
$exp(\frac{1}{2}(cos(cx)+cos(xy)))+5$
$50$
$100$
Perturbed Source - 3
$\frac{1}{4}(-a*exp(-b(\sqrt{\frac{1}{2}(x^{2}-y^{2})})))-\frac{3}{4}(exp(\frac{1}{2}(cos(cx)+cos(cy))))+a+exp(1)$
$50$
$100$

### v3

@@ 6 Conclusion
AI and ML frameworks have had a strong impact in the science and engineering communities. On the other hand, with the ever-rising application of data-driven surrogates build on multiple sources of data that it is gathered either experimentally or computationally, not much attention has been paid to the quality, completeness, and comprehensiveness of the excluded and unknown underlying parameters involved in the data generation process. These parameters vary from differences in manufacturing process, testing technique, and more. Through the research done in this paper, an interpretable multi-source data fusion framework based on Latent Variable Gaussian Process (LVGP) is proposed and extensively studied. Through LVGP models, a categorical variable was introduced to represent sources of information during model training and predictions. A low-dimensional latent space is then learned to incorporate the categorical source variable into multi-source data fusion. Through latent variable representation, both the excluded and underlying unknown physical parameter variations among the different sources are captured and introduced into the modelling, making the multi-source ML model interpretable and aware of the differences and similarities between the information sources. The inclusion of the data sources not only leads to more accurate models but also enables the study of source specific response surfaces, material optimization, and sensitivity analysis. Compared to using single source data for ML, the proposed data fusion method using multi-source information can provide a better prediction for sparse-data problems by taking advantage of the information of correlations and relationships among different data sources. With the quantified uncertainty from LVGP for the specified information source, the approach can further guide adaptive sampling or active learning when allocating resources for data collection. Additionally, a novel dissimilarity metric is also defined through the latent variables of the LVGP models that can help understand the aggregated differences and similarities in the information sources. The proposed approach is applied on and analyzed through two mathematical (representative parabola and 2D Ackley function) and two materials design (design of FeCrAl alloys and design of SmCoFe alloys) case studies. Finally, a novel application of the framework would involve performing materials design and development through autonomous experimentation studies where multiple sources of experimental data are gathered to guide the study. It is envisioned that the proposed multi-source framework is not limited to experimental or computational information sources and can be further extended to mixture of experimental and computational sources in numerous engineering and science applications.
@@ Acknowledgments
Northwestern faculty and student are grateful for the sponsorship from NSF 2219489 and the Army Research Laboratory under Cooperative Agreement Number W911NF-22-0121. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work is supported by the US Department of Energy, National Nuclear Security Administration, under award number DE-NE0009047. This material is based upon work supported by the U.S. Department of Energy’s Office of Energy Efficiency and Renewable Energy (EERE) under the Advanced Manufacturing Office, Award Number DE-AC0206H11357. The views expressed herein do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
[…]
@@
Information Source Function Formulation $(a=20,b=0.2,c=2)$ Number of Training Samples Number of Testing Samples
Ground Source $z=-a*exp(-b(\sqrt{\frac{1}{2}(x^{2}-y^{2})}))-exp(\frac{1}{2}(cos(cx)+cos(cy)))+a+exp(1)$ $20$ $100$
Perturbed Source - 1 $z=-a*exp(-b(\sqrt{\frac{1}{2}(x^{2}-y^{2})}))+10$ $50$ $100$
Perturbed Source - 2 $z=exp(\frac{1}{2}(cos(cx)+cos(xy)))+5$ $50$ $100$
Perturbed Source - 3 $z=\frac{1}{4}(-a*exp(-b(\sqrt{\frac{1}{2}(x^{2}-y^{2})})))-\frac{3}{4}(exp(\frac{1}{2}(cos(cx)+cos(cy))))+a+exp(1)$ $50$ $100$